Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by chaitanyareddy » Sat Aug 21, 2010 9:23 am
A and B alternately toss a coin. The first one to turn up a head wins. if no more
than five tosses each are allowed for a single game.
1- Find the probability that the person who tosses first will win the game?
2- What are the odds against A's losing if she goes first?

Please explain the solution for this.
Join the discussion
Source: — Problem Solving |

by Stuart@KaplanGMAT » Sat Aug 21, 2010 10:23 am
chaitanyareddy wrote:A and B alternately toss a coin. The first one to turn up a head wins. if no more
than five tosses each are allowed for a single game.
1- Find the probability that the person who tosses first will win the game?
2- What are the odds against A's losing if she goes first?

Please explain the solution for this.
Probability = (# of desired outcomes)/(total # of possibilities)

Fairly simple game, so let's brute force it by writing out strings of possible results:

1) H

game over, person 1 wins

2) TH

game over, person 2 wins

3) TTH

game over, person 1 wins

4) TTTH

game over, person 2 wins

5) TTTTH

game over, person 1 wins

6) TTTTT

game over, no winner

Question 1: probability that 1st person wins

# of scenarios in which she wins: 3
Total # of scenarios: 6

Prob(person 1 wins) = 3/6 = 1/2

Question 2: probably that 1st person doesn't lose

# of scenarios in which she doesn't lose: 4 (tie isn't a loss)
Total # of scenarios: 6

Prob(person 1 doesn't lose) = 4/6 = 2/3

Another way to look at question 2:

"Odds against her losing" = 1 - (odds of her losing)

# of scenarios in which she loses: 2
Total # of scenarios: 6

Odds of her losing = 2/6 = 1/3

Odds against her losing = 1 - 1/3 = 2/3
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by hala_absi » Sun Aug 22, 2010 6:58 am
Hello Stuart,
I want to ask you, I see that you assumed that the total number of tosses all together is 5 while it should be that "EACH" has a maximum of 5 tosses.
So if we say we have person A & B

each has either an H or T to get

Toss 1)
A:H B:T A wins
A:T B:H B wins
A:T B:T no one wins
Toss 2) (reached because no one won)
A:TH B:TT A wins
A:TT B:TH B wins
A:TT B:TT no one wins
Toss 3) (reached because no one won)
A:TTH B:TTT A wins
A:TTT B:TTH B wins
A:TTT B:TTT no one wins
Toss 4) (reached because no one won)
A:TTTH B:TTTT A wins
A:TTTT B:TTTH B wins
A:TTTT B:TTTT no one wins
Toss 5) (reached because no one won)
A:TTTTH B:TTTTT A wins
A:TTTTT B:TTTTH B wins
A:TTTTT B:TTTTT no one wins


So the total of A winning is 5, total B winning is 5, total no one winning is 5
Total scenarios are 15.
Probability of A winning is 5/15 = 1/3

Probability of A not losing is either winning or a tie = 10/15 = 2/3
Odds against her losing = 1- odds of her losing = 1- 5/15 = 1-1/3 = 2/3

Isnt this the right way of getting the answer?
Best Regards,
Hala Absi
Senior Software Engineer
Join the discussion

by gmat1011 » Sun Aug 22, 2010 8:12 am
For 1, why can't we consider it this way? why is this wrong:

- first tosser tosses a head on 1st toss and wins
- or, first tosser wins on his second toss
- or, first tosser wins on his third toss
- or, first tosser wins on his fourth toss
- or, first tosser wins on his fifth toss

each of the above requires the second tosser to toss a tail when he gets a chance to toss

prob = 341/512
Join the discussion